I am having a lot of trouble with this word problem in my Advanced 12th Grade Algebra class. I've studied back in the book, asked my parents to help, and worked it over and over, and I just can't get it. If someone could help me as soon as possible, I would really appreciate it! Thanks! :)
"The total curb weight5 of a Toyota Tundra truck, a Ford Mustang car, and a Smart For Two car is 11,150 lb. The weight of a Mustang is 5 lb less than the weight of two Smart For Two cars. The Tundra weighs 2135 lb more than the Mustang. What is the curb weight of each vehicle?"

Answer by graphmatics(170)

Let x be the weight of the Smart Car, then 2x-5 is the weight of the Mustang and 2x-5+2135 is the weight of the Tundra. We are given that:
x+2x-5+2x-5+2135=11150
5x=11150+10-2135
x=9025/5=1805 This is the Smart Car weight.
Tundra weight 2x-5=3605 .
Mustang weight 3605+2135=5740.